OpenStudy (anonymous):

2x2 – 4x = 0 A. {0, 2} B. {0, –2} C. {–6, –2} D. {0, 1/2}

OpenStudy (anonymous):

Hey is that 2x with two as an exponent or is it 2x times 2? The beginning is confusing

OpenStudy (anonymous):

its squared 2x2 is squared

OpenStudy (anonymous):

Ohh ok :) lemme help if I can

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

I think it's A

OpenStudy (anonymous):

do you have proof?

OpenStudy (chrisdbest):

I know it's A

Nnesha (nnesha):

what is the common factor ?

OpenStudy (anonymous):

When u do the math it's obvious that ud get 0 and 2 is what will be left over it all so I just put{ 0,2}

OpenStudy (anonymous):

ok hld on 1 sec

OpenStudy (chrisdbest):

Plug in 2 for x and you will get 2(2^2) - 4(2) = 0

OpenStudy (chrisdbest):

So 2(2^2) = 8, and 4(2) = 8. 8 - 8 = 0!!

OpenStudy (chrisdbest):

What!?? No medal?

OpenStudy (anonymous):

Thank you for the medal :)